# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@
Abstract
For a noncommutative algebra $\mathcal{A}$ and an antilinear automorphism $\rho$ of $\mathcal{A}$ there is a notion of a positive trace. Positive traces appear in study of supersymmetric gauge theories. We classify positive traces on $\mathcal{A}$ in two cases. The first case is when $\mathcal{A}$ is a quantization of a Kleinian singularity of type $D$, with certain restriction on the quantization parameter. The second case is when $\mathcal{A}=K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$ is an algebra containing $K$-theoretic Coulomb branches of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories.
0.1 Introduction
@@
Let $\rho$ be an antilinear automorphism of $\mathcal{A}$ and $g=\rho^{2}$. A $g$-twisted trace $T$ is called positive if $T(a\rho(a))>0$ for all nonzero $a\in\mathcal{A}$. In this case $(a,b)=T(a\rho(b))$ is a positive definite Hermitian form on $\mathcal{A}$ satisfying $(a\rho(b),c)=(a,bc)$.
@@
In the case when $\mathcal{A}$ is filtered, Etingof and Stryker ES (20) proved that twisted traces satisfying a certain nondegeneracy condition correspond to short nondegenerate star-products on $\operatorname{gr}\mathcal{A}$ introduced by Beem, Peelaers and Rastelli BPR (16). Positive traces correspond to unitary short star-products. In the case when $\mathcal{A}$ comes from a superconformal field theory, there is a certain unitary short star-product on $\operatorname{gr}\mathcal{A}$ that reflects the properties of the corresponding field theory. In some cases, a positive trace is unique up to scaling, which allows us to compute the corresponding unitary short star-product.
Another physical motivation, which does not require $\mathcal{A}$ to be filtered, comes from the work of Gaiotto Gai (25) and Gaiotto and Teschner GT (24). When $\mathcal{A}$ is a quantized Higgs/Coulomb branch of a superconformal field theory, there is a certain physically motivated choice of a conjugation $\rho$ and a trace $T$. It is expected that $T$ is positive. If this is the case, let $\mathcal{H}$ be a completion of $\mathcal{A}$ with respect to (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Then $\mathcal{A}$ acts on $\mathcal{H}$ by unbounded operators of left and right multiplication $L_{a}$, $R_{a}$. It is expected that we can choose the domain of the operators so that $L_{a}^{*}=R_{\rho(a)}$. Also, the image $v_{0}$ of $1\in\mathcal{A}$ in $\mathcal{H}$ satisfies $L_{a}v_{0}=R_{a}v_{0}$. The algebra $\mathcal{A}$ with the action on $\mathcal{H}$ provides more information about the field theory than the algebra $\mathcal{A}$ on its own.
@@
In this article we classify positive traces in the following two cases, described in more details below. First is when $\mathcal{A}$ is a deformation of a Kleinian singularity of type $D$ satisfying a certain condition. Second is when $\mathcal{A}$ is an algebra containing quantized $K$-theoretic Coulomb branches of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories.
@@
A Kleinian singularity is a quotient $\mathbb{C}^{2}/G$, where $G$ is a finite subgroup of the special linear group $\operatorname{SL}_{2}(\mathbb{C})$. The Kleinian singularity $\mathbb{C}^{2}/G$ has as its coordinate ring the commutative graded algebra $A=\mathbb{C}[u,v]^{G}$ of $G$-invariant polynomials in two variables.
@@
$\begin{pmatrix}e^{\frac{2\pi i}{n}}&0\\ 0&e^{-\frac{2\pi i}{n}}\end{pmatrix}.$
@@
$\begin{pmatrix}e^{\frac{\pi i}{n}}&0\\ 0&e^{-\frac{\pi i}{n}}\end{pmatrix}\quad\text{and}\quad h=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},$
@@
which contains $\mathbb{Z}/2n\mathbb{Z}$ as a subgroup of order two. The element $h$ defines an automorphism $\mathbb{C}[u,v]^{Z/2n\mathbb{Z}}$ of order two and the invariant subalgebra is precisely the Kleinian singularity of type $D_{n}$.
@@
Kleinian singularities are graded algebras. We can look at their noncommutative filtered deformations, that is, filtered associative algebras $\mathcal{A}$ with associated graded algebra $\operatorname{gr}\mathcal{A}$ equal to $A$, and try to classify positive traces on these deformations. Twisted and positive traces on deformations and $q$-deformations of Kleinian singularities of type $A$ were classified by the first author with Etingof, Rains, Stryker in EKRS (21) and in Kly (22). From a physical point of view positive traces were studied in DG (21); DFPY (18); DPY (16).
@@
In the first chapter, we consider the next case, deformations of type $D$. Not all deformations of Kleinian singularities of type $D$ can be obtained as a subalgebra inside a Kleinian singularity of type $A$, but in the first chapter we consider the ones that do. Moreover, we assume that the quantization parameter is generic.
@@
We obtain classification results for traces on deformations of Kleinian singularities of type D analogous to those obtained in EKRS (21) for type A. Most importantly, we show that, under certain assumptions, all positive traces of filtered deformations of Kleinian singularities of type D are restrictions of positive traces of filtered deformations of Kleinian singularities of type A.
@@
We begin by outlining the technical preliminaries for our work in Section 1.1. Then, in Section 1.2, we introduce several algebra elements which are important for our results and prove various identities and relations among them. We characterize traces of deformations of Kleinian singularities of type D that satisfy the assumptions above in Section 1.3, and use this characterization to obtain an analytic formula for traces in Section 1.4. Finally, in Section 1.5, we prove our main result regarding positive traces.
We now turn to the second chapter.
@@
Let $G$ be a complex reductive group and $N$ be a representation of $G$. One can construct a supersymmetric gauge theory from this data. There are two algebraic varieties associated to this theory—the Higgs and Coulomb branches. A mathematical definition of Coulomb branches was given by Braverman, Finkelberg and Nakajima in BFN (18)
Each Coulomb branch comes with a quantization of its ring of functions. Quantized Coulomb branches include well-known algebras such as spherical double affine Hecke algebras or central reductions of $U(\mathfrak{sl}_{n})$
@@
In Lemma 6.9 of BFN (18) it is proved that in the case of $G=\operatorname{SL}_{2}(\mathbb{C})$ Coulomb branches are exactly type $D_{n}$ Kleinian singularities $xy^{2}=z^{2}+x^{n-1}$ plus three exceptions: the cases $n=2$, $n=1$ of this equation, and the surface $xy^{2}=z^{2}+y$. If we take representation $N=(\mathbb{C}^{2})^{m}$, we get $n=m$. (in the notation of BFN (18), the representation is denoted by $\mathbf{N}$ and our $n$ corresponds to $N$ in Lemma 6.9).
This fact motivates the title of the article, but we do not use any facts about Coulomb branches in the first chapter.
@@
Coulomb branches are defined as homology of an infinite-dimensional scheme. Changing homology to K-theory we also get an algebra, called K-theoretic Coulomb branch algebra. Physically, we take 4-dimensional gauge theory with the same group $G$ and matter $M$, and compactify it on a circle. Algebraically, a deformed $K$-theoretic Coulomb branch is a certain $q$-deformation of a usual Coulomb branch.
In the second part of the paper we consider the algebra $\mathcal{A}$ introduced by Gaiotto and Teschner in Section 3.5 of GT (24). This algebra contains $K$-theoretic Coulomb branches of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories, where pure means that $N=0$, and in this case the BFN construction gives the $G(O)$-equivariant $K$-theory of an affine Grassmannian $\operatorname{Gr}_{G}$.
Let $v,w$ be variables satisfying $wv=qvw$; they generate a q-Weyl algebra. We consider the localized q-Weyl algebra $W$, inverting $v,w$ and $q^{k}v-q^{-k}v^{-1}$ for all $k\in\mathbb{Z}$.
Then $\mathcal{A}$ as a subalgebra of $W$ generated by $v^{n}+v^{n}$ and $H_{a}=q^{\frac{a}{2}}(v-v^{-1})^{-1}(v^{a}w+v^{-a}w^{-1})$. This description is equivalent to the one in Section 3.5.1 of GT (24) by taking $u_{+}=(v-v^{-1})^{-1}w$ and $u_{-}=(v-v^{-1})^{-1}w^{-1}$.
The algebra $\mathcal{A}$ should be equal to $K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})$, an object similar to the $K$-theoretic Coulomb branch of pure $\operatorname{PGL}(2)$ gauge theory, but the $K$-theory is taken with respect to $\operatorname{SL}(2)$, not $\operatorname{PGL}(2)$. More precisely, we consider an algebra generated by the images of dressed miniscule monopole operators and K-theory of a point under the localization map $(i_{*})^{-1}\colon K^{\operatorname{SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}(\operatorname{Gr}_{\operatorname{PGL}_{2}})\to K^{\mathbb{C}^{\times}_{v}\times\mathbb{C}^{\times}_{q}}(\operatorname{Gr}_{T})_{loc}$, where the multiplicative subset on the right is defined so that $W=K_{\mathbb{C}^{\times}_{v}\times\mathbb{C}^{\times}_{q}}(\operatorname{Gr}_{T})_{loc}$. It can be checked that the image of the class corresponding to the sheaf $O(l)$ (with trivial loop torus, $\mathbb{C}^{\times}_{q}$, action) on a smooth $G(\mathbb{C}[[t]])$-orbit $\mathbb{P}^{1}\subset\operatorname{Gr}_{\operatorname{PGL}_{2}}$ in $W$ is $-(v-v^{-1})^{-1}(v^{l+1}w-v^{-l-1}w^{-1})$, so that the result is precisely the algebra $\mathcal{A}$.
[…]
@@
While $\mathcal{A}$ does not exactly fit into the sequence of $q$-deformations of Kleinian singularities of type $D$, one could use a similar strategy to classify positive traces on deformed $K$-theoretic Coulomb branches with gauge group $\operatorname{SL}(2)$ or $\operatorname{PGL}(2)$. Also, Kauffman skein algebras should be equal to the “mixed” Coulomb branches, where, as above, the group for the space of triples and the group in equivariant K-theory are different quotients of the same simply-connected Lie group. For adjoint representation, the flavor deformations of $K^{\operatorname{SL}(2,\mathbb{C}[t])\rtimes\mathbb{C}_{q}^{\times}}(\mathcal{R}_{\operatorname{PGL}(2),\mathfrak{sl}_{2}})$ correspond to the skein algebras for one-punctured torus; this is proved by Allegretti and Shan AS (24) in the paragraph after Proposition 5.5.
[…]
The first step in classification is to describe the full image of $\mathcal{A}$ in $W$. Compared to the above, we use slightly different inclusion, given by right fractions and not left, in the second chapter. In Section 2.1 we prove the following
Proposition (2.1.1).
The algebra $\mathcal{A}$ coincides with the set of elements $a=\sum w^{i}f_{i}$ of $W$ such that all $f_{i}$ have simple poles only possibly at ±$\pm q^{j}$, $j\in\mathbb{Z}$ and the corresponding residues $f_{i}^{(j,\pm)}$ satisfy
  1. 1.
    $f_{-i}(v^{-1})=(-1)^{i}f_{i}(v)$;
  2. 2.
    $f_{i}^{(i,\pm)}=0$;
  3. 3.
    For a positive integer $k$ we have $f_{i+k}^{(i,\pm)}=(-1)^{k+1}f_{i-k}^{(i,\pm)}$, where $\pm$ is the same sign on both sides.
The algebra $\mathcal{A}$ has a $\mathbb{Z}/2$-grading by the degree of $w$. In Section 2.2 we prove that any twsited trace $T$ can be decomposed as $T=T_{0}+T_{1}$, where $T_{i}$ is supported on the $i$-th graded piece and write an analytic formula for $T_{0}$.
In Section 2.3 we prove that for a positive trace $T$ we should have $T_{1}=0$ and $T_{0}(\sum w^{i}a_{i})=\int_{S^{1}}a_{0}(z)\omega(z)\frac{dz}{z}$, where $\omega(z)$ is a quasi-periodic function vanishing at ±1$\pm$ 1± 1.
Then we express the positivity condition on $T$ in terms of the behavior of $w$ on $S^{1}$ and $\sqrt{q}S^{1}$. Using this, we can compute the dimension of the cone of possible traces. In particular, for the antilinear automorphism $H_{a}\mapsto H_{a+2}$, there exists a unique positive trace. Taking into account the difference between the conventions in physics and mathematics, this corresponds to the automorphism considered by Gaiotto and Teschner in Section 3.5 of GT (24).
0.2 Acknowledgements
The first chapter was written by the J.V. under the mentorship of D.K. at Research Science Institute, hosted by the Massachusetts Institute of Technology. We would like to thank Pavel Etingof for suggesting this project and for comments on the earlier version of the paper. We would also like to thank RSI Head Mentor, Dr. Tanya Khovanova for comments on the earlier versions of the paper.
J.V. also thanks Peter Gaydarov for helpful suggestions in preparing this paper, and Allen Lin and Prof. John Rickert for their thoughtful comments on an earlier version of this paper.
D.K. would like to thank Davide Gaiotto for suggesting the problem discussed in the second chapter and to Artem Kalmykov for providing the SAGE code for computing multiplication in the algebra $\mathcal{A}$. D.K. started working on the article when he was a graduate student and then a summer research specialist at MIT.
Chapter 1 Deformations of type $D$
1.1 Preliminaries

### v2

# Positive Traces on Certain ${\rm SL}(2)$ Coulomb Branches

@@
[…]
Positive Traces on Certain ${\rm SL}(2)$ Coulomb Branches
[…]
Positive Traces on Certain $\boldsymbol{{\rm SL}(2)}$ Coulomb Branches
[…]
Daniil KLYUEV a and Joseph VULAKH b
[…]
D. Klyuev and J. Vulakh
[…]
Received July 28, 2025, in final form March 15, 2026; Published online April 10, 2026
[…]
For a noncommutative algebra $\mathcal{A}$ and an antilinear automorphism $\rho$ of $\mathcal{A}$, there is a notion of a positive trace. When we have a three-dimensional $\mathcal{N}=4$ gauge theory or four-dimensional $\mathcal{N}=2$ gauge theory compactified on a circle, classification of positive traces on its Coulomb branch $\mathcal{A}$ can give a better understanding of this theory. We classify positive traces on $\mathcal{A}$ in two cases. The first case is when $\mathcal{A}$ is a quantization of a Kleinian singularity of type $D$, with certain restriction on the quantization parameter. The second case is when $\mathcal{A}=K^{{\rm SL}(2,\mathbb{C}[[t]])\rtimes\mathbb{C}_{q}^{\times}}({\rm Gr}_{{\rm PGL}_{2}})$ is an algebra containing $K$-theoretic Coulomb branches of pure ${\rm SL}(2)$ and ${\rm PGL}(2)$ gauge theories.
[…]
Coulomb branches; twisted traces; gauge theory; sphere quantization; Schur quantization
[…]
53D50; 81R10; 14D21
1 Introduction
@@
Let $\rho$ be an antilinear automorphism of $\mathcal{A}$ and $g=\rho^{2}$. A $g$-twisted trace $T$ is called positive if $T(a\rho(a))>0$} for all nonzero $a\in\mathcal{A}$. In this case, $(a,b)=T(a\rho(b))$ is a positive definite Hermitian form on $\mathcal{A}$ satisfying $(a\rho(b),c)=(a,bc)$.
@@
In the case when $\mathcal{A}$ is filtered, Etingof and Stryker [12] proved that twisted traces satisfying a certain nondegeneracy condition correspond to short nondegenerate star-products on $\operatorname{gr}\mathcal{A}$ introduced by Beem, Peelaers and Rastelli [2]. Positive traces correspond to unitary short star-products. In the case when $\mathcal{A}$ comes from a superconformal field theory, there is a certain unitary short star-product on $\operatorname{gr}\mathcal{A}$ that reflects the properties of the corresponding field theory. In some cases, a positive trace is unique up to scaling, which allows us to compute the corresponding unitary short star-product.
Another physical motivation comes from the work of Gaiotto [13] and Gaiotto and Teschner [15]. When $\mathcal{A}$ is a quantized Higgs/Coulomb branch of a superconformal field theory, there is a certain physically motivated choice of a conjugation $\rho_{\text{sph}}$ and a trace $T_{\text{sph}}$. We call $T_{\text{sph}}$ sphere trace. Physicists expect $T_{\text{sph}}$ to be positive and it would be useful to have a mathematical proof of that. The sphere trace was introduced by Gaiotto and Okazaki [14], where they proposed a formula for the sphere correlation functions of Coulomb/Higgs operators in terms of the sphere trace.
The algebra $\mathcal{A}$ does not have to be filtered, and Gaiotto and Teschner consider $K$-theoretic Coulomb branches. An advantage of $K$-theoretic Coulomb branches is that $\rho_{\text{sph}}$-positive trace should always be unique up to scaling. A mathematical proof of that is much more complicated than just proving positivity of $T_{\text{sph}}$ and the uniqueness statement is very useful for the following reason.
Suppose that $\mathcal{A}$ indeed has a unique $\rho_{\text{shp}}$-positive trace $T_{\text{sph}}$. In this case, we canonically produce a Hilbert space $\mathcal{H}$ with an action of $\mathcal{A}\times\mathcal{A}^{{\rm op}}$ as follows. The space $\mathcal{H}$ be a completion of $\mathcal{A}$ with respect to $(a,b)=T_{\text{sph}}(a\rho_{\text{sph}}(b))$ and $\mathcal{A}$ acts on $\mathcal{H}$ by unbounded operators of left and right multiplication $L_{a}$, $R_{a}$. The pair $(\mathcal{A},\mathcal{H})$ is useful for understanding the superconformal field theory that $\mathcal{A}$ came from. See [15, Section 1.3] for the discussion of class $\mathcal{S}$ theories and the uniqueness of the trace. The same motivation applies to $K$-theoretic Coulomb branches of not necessarily conformal supersymmetric gauge theories; see [15, Section 2.2] for the discussion of how do non-conformal theories fit into their picture.
@@
In this article, we classify positive traces in the following two cases, described in more details below. The first is when $\mathcal{A}$ is a deformation of a Kleinian singularity of type $D$ satisfying a certain condition; we study this case in Sections 26. The second is when $\mathcal{A}$ is an algebra containing quantized $K$-theoretic Coulomb branches of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories; we study this case in Sections 79. In particular, our results imply that the sphere trace $T_{\text{sph}}$ is positive when $\mathcal{A}$ is the quantized $K$-theoretic Coulomb branch of pure $\operatorname{SL}(2)$ or $\operatorname{PGL}(2)$ gauge theory.
@@
The first case we study deals with Kleinian singularities. A Kleinian singularity is a quotient $\mathbb{C}^{2}/G$, where $G$ is a finite subgroup of the special linear group $\operatorname{SL}_{2}(\mathbb{C})$. The Kleinian singularity $\mathbb{C}^{2}/G$ has as its coordinate ring the commutative graded algebra $A=\mathbb{C}[u,v]^{G}$ of $G$-invariant polynomials in two variables.
@@
$\begin{pmatrix}{\rm e}^{\frac{2\pi{\rm i}}{n}}&0\\ 0&{\rm e}^{-\frac{2\pi{\rm i}}{n}}\end{pmatrix}$.
@@
$\begin{pmatrix}{\rm e}^{\frac{\pi{\rm i}}{n}}&0\\ 0&{\rm e}^{-\frac{\pi{\rm i}}{n}}\end{pmatrix}\qquad\text{and}\qquad h=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$,
@@
which contains $\mathbb{Z}/2n\mathbb{Z}$ as a subgroup of index two. The element $h$ defines an automorphism of $\mathbb{C}[u,v]^{Z/2n\mathbb{Z}}$ of order two and the invariant subalgebra is precisely the Kleinian singularity of type $D_{n-1}$. When the value of $n$ is not relevant, we refer to such singularities as being of type D without reference to $n$.
@@
Kleinian singularities are graded algebras. We can look at their noncommutative filtered deformations, that is, filtered associative algebras $\mathcal{A}$ with associated graded algebra $\operatorname{gr}\mathcal{A}$ equal to $A$, and try to classify positive traces on these deformations. Twisted and positive traces on deformations and $q$-deformations of Kleinian singularities of type $A$ were classified by the first author with Etingof, Rains, Stryker in [11] and in [17]. From a physical point of view, positive traces were studied in [7, 8, 9, 10].
@@
In the first part of the article, we consider the case of deformations of type $D$. Not all deformations of Kleinian singularities of type $D$ can be obtained as a subalgebra inside a Kleinian singularity of type $A$, but in the first part we consider the ones that do. Moreover, we assume that the quantization parameter is generic.
@@
We obtain classification results for traces on deformations of Kleinian singularities of type D analogous to those obtained in [11] for type A. Our most important result in this direction is the following theorem.
{theorem*}
[Theorem 6.5] Under assumptions described in Section 6, all positive traces of filtered deformations of Kleinian singularities of type D are restrictions of positive traces of filtered deformations of Kleinian singularities of type A.
@@
We outline the technical preliminaries for our work on deformations of Kleinian singularities in Section 2. Then, in Section 3, we introduce several algebra elements which are important for our results and prove various identities and relations among them. We characterize traces of deformations of Kleinian singularities of type D that satisfy the assumptions above in Section 4, and use this characterization to obtain an analytic formula for traces in Section 5. Finally, in Section 6, we prove our main result regarding positive traces of filtered deformations of Kleinian singularities of type D.
We now turn to the second part of the article.
@@
Let $G$ be a complex reductive group and $N$ be a representation of $G$. One can construct a supersymmetric gauge theory from this data. There are two algebraic varieties associated to this theory – the Higgs and Coulomb branches. A mathematical definition of Coulomb branches was given by Braverman, Finkelberg and Nakajima in [4].
In the case when $N=0$, the Coulomb branch was computed earlier in [3], it is isomorphic to the universal centralizer scheme $\mathfrak{Z}_{\check{\mathfrak{g}}}^{\check{G}}$ for the Langlands dual group. Other examples include slices in the affine Grassmannian for type $ADE$ simple Lie groups [5], in particular the nilpotent cone $\mathcal{N}_{\mathfrak{sl}_{n}}$ is a Coulomb branch. For the proof of the last assertion, see the last remark in [16], for example.
Each Coulomb branch $\mathcal{M}_{G,N}=\operatorname{Spec}\mathcal{A}_{G,N}$ comes with a quantization of its ring of functions $\mathcal{A}_{G,N}^{\hbar}$. For example, any quantization of the nilpotent cone algebra $\mathbb{C}[\mathcal{N}_{\mathfrak{sl}_{n}}]$ is a central reduction of $U(\mathfrak{sl}_{n})$. This example provides a link between positive traces on quantized Coulomb branches and infinite-dimensional unitary representations of simple complex Lie groups. For an example of how this works in the case when $n=2$, see [18, Section 4.3, pages 43 and 46].
@@
In [4, Lemma 6.9], it is proved that in the case of $G=\operatorname{SL}_{2}(\mathbb{C})$ Coulomb branches are exactly type $D_{n}$ Kleinian singularities $xy^{2}=z^{2}+x^{n-1}$ plus three exceptions: the cases $n=2$, $n=1$ of this equation, and the surface $xy^{2}=z^{2}+y$. If we take the representation $N=\bigl(\mathbb{C}^{2}\bigr)^{m}$, we get n=m{n=m}. In the notation of [4], the representation is denoted by $\mathbf{N}$ and our $n$ corresponds to $N$ in Lemma 6.9. This fact motivates the title of the article, but we do not use any facts about Coulomb branches in the first part of the article.
@@
The Coulomb branch algebra $\mathcal{A}_{G,N}$ is defined as equivariant homology of an infinite-dimensional scheme $\mathcal{R}_{G,N}$. Changing homology to K-theory, we also get an algebra, called the K-theoretic Coulomb branch algebra. Physically, we take 4-dimensional gauge theory with the same group $G$ and matter $M$, and compactify it on a circle. Algebraically, a deformed $K$-theoretic Coulomb branch algebra $\mathcal{A}_{G,N}^{q}$ is a certain $q$-deformation of $\mathcal{A}_{G,N}$.
In the second part of the paper, we consider the algebra $\mathcal{A}$ introduced by Gaiotto and Teschner in [15, Section 3.5]. This algebra contains $K$-theoretic Coulomb branches of pure $\operatorname{SL}(2)$ and $\operatorname{PGL}(2)$ gauge theories in the following sense. Gaiotto and Teschner use the description of Coulomb branch via the abelianization map, also known as the localization map in the mathematical literature. In our case, the target of the localization map is a certain localization $W$ of a $q$-Weyl algebra $W_{0}=\mathbb{C}\langle u,v\rangle/(vw-qwv)$. Then $\mathcal{A}$ is generated by the images of $\mathcal{A}_{\operatorname{SL}(2),0}^{q}$ and $\mathcal{A}_{\operatorname{PGL}(2),0}^{q}$ in $W$. In Section 7, we will describe $\mathcal{A}$ and its connection to the BFN construction and show that there exist involutions $\tau_{1}$, $\tau_{2}$ of $\mathcal{A}$ such that $\mathcal{A}_{SL(2),0}^{q}=\mathcal{A}^{\tau_{1}}$ and $\mathcal{A}_{\operatorname{PGL}(2),0}^{q}=\mathcal{A}^{\tau_{2}}$, the fixed point subalgebras.
@@
While $\mathcal{A}$ does not exactly fit into the sequence of $q$-deformations of Kleinian singularities of type $D$, one could use a similar strategy to classify positive traces on deformed $K$-theoretic Coulomb branches with the gauge group $\operatorname{SL}(2)$ or $\operatorname{PGL}(2)$. Also, Kauffman skein algebras should be equal to the “enlarged” Coulomb branches constructed similarly to Section 7. Allegretti and Shan proved in [1] in the paragraph after Proposition 5.5 that the skein algebra for the once-punctured torus is isomorphic to $K^{\operatorname{SL}(2,\mathbb{C}[t]\times\mathbb{C}[t]^{\times})\rtimes\mathbb{C}_{q}^{\times}}(\mathcal{R}_{\operatorname{PGL}(2),\mathfrak{sl}_{2}})$. The latter algebra is the universal flavor deformation of a modified BFN construction in the case when $N=\mathfrak{sl}_{2}$ is the adjoint representation.
We classify positive traces on $\mathcal{A}$ with respect to an involution $\rho=\rho_{m}$ that depends on an integer $m$. In the case when $m=2$ this involution coincides with the one considered by Gaiotto and Teschner, and we further prove that the positive trace is unique up to scaling.
In Section 7, we describe $\mathcal{A}$ and prove a proposition that describes the full image of $\mathcal{A}$ in the localized $q$-Weyl algebra $W$. In Section 8, we describe traces in terms of a quasi-periodic function $\omega$ holomorphic on $\mathbb{C}^{\times}$. In Section 9, we express the positivity condition on $T$ in terms of the behavior of $\omega$ on $S^{1}$ and $\sqrt{q}S^{1}$, finishing the classification. The answer is given by the following theorem.
{theorem*}
[Theorem 9.4] Positive traces on $\mathcal{A}$ are in one-to-one correspondence with holomorphic functions $\omega$ on $\mathbb{C}^{\times}$ such that
  1. $(1)$
    $\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z)$;
  2. $(2)$
    $\omega(z)=\omega\bigl(z^{-1}\bigr)$;
  3. $(3)$
    $\omega(1)=\omega(-1)=0$;
  4. $(4)$
    $\omega(z)$ and $z^{-\frac{m}{2}}\omega\bigl(\sqrt{q^{-1}}z\bigr)$ are nonnegative on $S^{1}$.
The convex cone of such functions $\omega$ has real dimension $\frac{m}{2}-1$. In particular, for $m=4$ there is a unique positive trace up to a constant.
Part 1: Deformations of type $\boldsymbol{D}$
2 Preliminaries